Bob’s Bakery sells three types of cupcakes, Chocolate with a berry on top, Vanilla with an icing face, Strawberry with sprinkles. The following table shows the sales price and variable costs for each type. The bakery incurs $300,000 a year in fixed expenses. Assume that it sells two Chocolate for every one Vanilla and every one Strawberry.
Cupcake type
Sales price
Variable cost
Chocolate with a berry on top
$0.35
$0.20
Vanilla with an icing face
$0.50
$0.45
Strawberry with sprinkles
$0.40
$0.27
Required
 How many cupcakes of each type will be sold at breakeven point?
 What amount of revenue would need to be generated by each type of cupcake for the company to earn $60,000 in operating income?
Answer #1
Particular chocolate vanilla Strawberry
Sales $0.35 $0.50 $0.40
Less variable cost $0.20 $0.45 $0.27
Contribution per unit $0.15 $0.05 $0.13
Sales mix 2 1 1
Total $0.30 $0.05 $0.13

Total contribution = $0.30 + $0.05 + $0.13 = $0.48

1. To break even units required = fixed cost / contribution = $300000/$0.48 = 625000

Hence unite required

Chocolate = 625000*2 = 1250000

Vanilla = 625000*1 = 625000

Strawberry = 625000*1 = 625000

2. Required profit = $60000

Units required = ($60000+$300000) / $0.48 = 7500000

Units required of

Chocolate = 750000*2 = 1500000

Vanilla = 750000*1 = 750000

Strawberry = 750000*1 = 750000

Feel free to ask any queries..

Also plz upvote it means a lot .. thank you